# Mamo: a Mathematical Modeling Benchmark with Solvers

### v1

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
@@
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solves, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of large language models (LLMs). Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
The contributions of this paper are as below:
  1. 1.
    Clarification of a Mathematical Modeling Framework for LLMs in Natural Language Domains: We have meticulously clarified the concept of mathematical modeling within Large Language Models (LLM), focusing on the realm of natural language challenges. This effort has culminated in the establishment of a comprehensive framework designed to guide the mathematical modeling process in LLMs, addressing both theoretical underpinnings and practical applications.
  2. 2.
    Innovation in Benchmarking Strategy through Solver Integration: Our research introduces a pioneering benchmarking strategy for the assessment of mathematical modeling process. By incorporating solvers into the benchmarking methodology, this strategy facilitates a rigorous evaluation process. This novel approach sets a new standard for the assessment in NLP research.
  3. 3.
    Creation of the Mamo Benchmark for Extensive Mathematical Modeling Assessment: We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions, focusing on areas such as ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks. With a total of 1059 meticulously curated questions, Mamo stands as a comprehensive tool for assessing the breadth and depth of mathematical modeling proficiency.
@@ 3.1 Mathematical Modeling
Motivations to benchmark mathematical modeling for LLMs
We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
@@ 4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$). This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
@@ 4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Tab. 1.
Category In Mamo
Changes and Differences
Proportionality and Geometric Similarity
Optimization Problem Modeling
Differential Equations
Probabilistic Modeling
Table 1: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
@@ 5.1 Data Selection and Source Credibility
The data set for our benchmark consists of a blend of manually selected and GPT-generated questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance.
Annotator Qualifications
The benchmarking process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines.  [^2] See the details of the annotator qualifications in Appendix. B:
5.1.1 Manually Selected Questions
@@
  • Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question.
  • Rewriting Process: Each question is meticulously rewritten to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology. This process also circumvents potential licensing issues, as the original content is significantly transformed.
Data Synthesis
@@
The process of annotating data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix. E.
Cross Validation
GPT-4 is employed to confirm that the newly generated questions are amenable to being solved by solvers using the final-state approach, ensuring the benchmark’s focus remains on the LLM’s modeling capabilities. Moreover, we employ a comprehensive review process, re-examining each question for consistency and correctness, thereby ensuring the quality of our dataset. See the review details in Appendix. A.
5.2 Data Statistics and Visualization
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels.
The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP.
See the word cloud of the combined data in Figure 2
5.3 Guidance of Data Synthesis
@@
The synthesis of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
@@
Final-State Approach Solvability: Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
@@
Numerical Answers: The answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
@@
Significant Figures and Precision: Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
@@
Modeling Capacity Test: To specifically test the LLM’s modeling capabilities:
  • For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$.
  • For optimization problems, questions should aim to find the optimal value.
These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem.
Property 5.
Real-World Problem Context: Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
@@ 6.1.1 Evaluation Details
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix. (see Appendix C and Appendix. G).
@@
It is essential to distinguish errors in modeling from those in coding or file formatting. Therefore, if the LLM’s output contains syntax errors or incorrect .lp formatting, we need to rectify these issues without altering the underlying logic of the code or model. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax errors, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the orignal LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifier in Section 6.3 and the prompts in Appendix.  G. Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like " “‘python “‘ " that will affect our testing.
@@ 6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  333 See the evaluation script in Appendix. F
Precision Adjustment for Floating-Point Numbers If the standard answer, $A$, is a floating-point number, the precision adjustment involves calculating the number of significant digits, $n$, after the decimal point. Both $O$ and $A$ are then scaled by $10^{n}$ for comparison.
General Comparison Criterion The overall correctness is determined based on the following conditions, applicable to both adjusted and original values:
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{ or }|O-A|\leq 1,\\ 0&\text{otherwise.}\end{cases}$ (1)
This structured evaluation ensures that the LLM’s output is accurately assessed against the standard answer with consideration for the specific requirements of numerical precision and the inherent variability in computational answers.
@@ 6.2 Benchmarking Results
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o 69.39% 73.64% 42.50% 67.63% 87.27% 23.22%
GPT-4o 69.90% 74.55% 42.50% 68.21% 87.42% 32.23%
GPT-4-turbo 66.84% 69.09% 42.50% 65.12% 87.88% 23.70%
GPT-4-turbo 67.86% 70.00% 42.50% 65.99% 87.88% 30.33%
GPT-4 60.71% 65.45% 42.50% 60.12% 86.50% 22.27%
GPT-4 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo 17.35% 16.36% 7.50% 15.90% 81.29% 9.95%
GPT-3.5-turbo 25.51% 31.82% 15.00% 26.30% 84.82% 9.95%
Claude-3-Sonnet 53.57% 61.82% 30.00% 53.47% 83.59% 25.12%
Claude-3-Sonnet 55.61% 63.64% 35.00% 55.78% 84.20% 25.59%
Claude-3-Haiku 42.35% 60.91% 20.00% 45.66% 78.53% 19.91%
Claude-3-Haiku 52.04% 63.64% 22.50% 52.31% 85.89% 21.33%
Gemini-1-pro 26.53% 35.45% 2.50% 26.59% 78.99% 15.31%
Gemini-1-pro 35.71% 39.09% 7.50% 33.53% 80.06% 15.31%
Gemini-1.5-pro 66.33% 66.36% 47.50% 64.16% 85.28% 30.81%
Gemini-1.5-pro 67.35% 66.36% 52.50% 65.32% 85.28% -
DeepSeek-v2 40.82% 45.45% 32.50% 41.33 % 86.50% 6.16%
DeepSeek-v2 42.35% 64.55% 32.50% 48.27% 86.66% 6.16%
DeepSeek-coder 32.14% 49.09% 15.00% 35.55% 75.77% 18.96%
DeepSeek-coder 34.18% 49.09% 15.00% 36.71% 76.99% 18.96%
Table 2: Evaluation result on Mamo Benchmark. The remark $\dagger$ indicates it additionally uses the original LLM to rectify syntax error of the code and .lp file generated by the corresponding LLM. Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09 and GPT-4o denotes GPT-4o-2024-05-13, while GPT-4 refers to GPT-4-0613

### v2

@@ 1 Introduction
Recent advancements in Large Language Models have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
@@
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solvers, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of LLMs. Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
The contributions of this paper are as below: (1) We introduce a new benchmarking strategy to assess mathematical modeling through Solvers, facilitating a rigorous evaluation process. (2) We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions (With a total of 1209 meticulously curated questions), including ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
@@ 3.1 Mathematical Modeling
Motivationsto Benchmark Mathematical Modeling for LLMs   We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
@@ 4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$), since typically the wrong ${\widehat{M}}$ will result in wrong ($\widehat{A}$), like the examples in Appendix  G. This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
@@ 4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Table 7 in Appendix F.
@@ 5.1 Data Selection and Source Credibility
The dataset for our benchmark consists of a blend of manually selected and GPT-synthetized questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance. The data collection and synthesis process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines  [^2] See the details of the annotator qualifications in Appendix C.
Manually Selected Questions
@@
The manually selected questions are high quality and reliable. (1)Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question. (2) Transformation: Each question is meticulously transformed to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
Question Formulation and Verification
@@
The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix K.
5.2 Data Quality Checking
In the ODE part, we employed GPT-4 to confirm the solvability of newly generated questions, filtering out invalid ones. A thorough evaluation identified questions lacking information, with incorrect answers, and unclear statements. Consequently, problematic questions were deleted or corrected, resulting in a refined dataset. In the optimization section, particularly with the easy LP part, invalid questions were filtered out, and those with incorrect answers or misleading descriptions were corrected. In the complex LP part, questions formulated with wrong mathematical models were corrected. The summary of the review process is presented in Table 1.
Category Initial Questions Filtered Final Questions
ODE Part 383 37 346
Easy LP Part 688 48 640
Complex LP Part 211 0 211
Table 1: Summary of the review process.
Reviewer Accuracy Rate
Reviewer 1 90.0%
Reviewer 2 84.0%
Reviewer 3 88.0%
Reviewer 4 74.0%
Table 2: Accuracy of reviewers.
For the cross-review process, 50 questions from the ODE dataset were selected for deeper analysis by four independent reviewers. Reviewers evaluated the questions, marking responses as either numeric answers or "error" if the question was problematic. The effectiveness of this review was quantified using metrics like Cohen’s Kappa to measure inter-rater reliability and accuracy rates for individual reviewers. The average Cohen’s Kappa was 0.60, indicating moderate to substantial agreement among reviewers, with individual Kappa values ranging from 0.50 to 0.71. Accuracy rates are moderately high as shown in Tables 2. These results highlight both the effectiveness of the review process and areas where alignment and training can improve consistency and accuracy among reviewers. See the review details in Appendix B.
5.3 Guidance of Data Collection
@@
The collection of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
@@
Final-State Approach Solvability  Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
@@
Unified and Numerical Answers: To handle potentially various valid solutions for one problem: For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$. For optimization problems, questions should aim to find the optimal value. These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem. Besides, the answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
@@
Significant Figures and Precision  Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
To assess the LLM’s capability to abstract mathematical models from real-world scenarios, we introduce Property 4.
@@
Real-World Problem Context  Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.
5.4 Data Statistics
In the section on Ordinary Differential Equations (ODE), we present a total of 346 problems: 196 are based on first-order equations, 110 on second-order equations, and 40 on systems of equations, offering a comprehensive exploration across different levels. The optimization section is divided into two segments: Easy LP, featuring 652 high school-level Mixed Integer Linear Programming (MILP) problems, and Complex LP, comprising 211 undergraduate-level problems that integrate both LP and MILP. See the word cloud of the combined data in the Appendix L.
@@ 6.1.1 Evaluation Details
In ODE parts, we ask LLMs to output Python code calling solvers ( solve_ivp odeint in SciPy, and dsolve in NumPy). In the optimization sections, we prompt LLMs to generate the standard .lp file. Compared to the testing methodology in NLP4LP (AhmadiTeshnizi et al., 2023b), which involves prompting the model to output Python code for invoking the solver, the .lp format more closely aligns with the standard optimization form, making it more readable. Furthermore, optimization solvers such as COPT (Ge et al., 2023) can directly read and solve the .lp file, minimizing potential issues related to the LLM’s coding capabilities. To better adhere to the format, we conducted tests using 3-shot learning. Detailed information about the number of few-shot samples used in our experiments and the result can be found in the Appendix (see Appendix I and Section O). Besides, at the beginning of the testing, we do the first clean of the codes, clearing up the string like ‘ “‘ python “‘ ’ that will affect our testing.
@@
It is essential to distinguish errors in modeling from those in coding or file formatting [^3] Appendix  H displays the example of formating errors instead of modeling errors.. Therefore, if the LLM’s output contains syntax errors in python code or .lp code, we need to rectify these issues without altering the underlying logic of the code. This step ensures that our evaluation focuses solely on the modeling process. To achieve this, we use a language model to fix any syntax error, referring to it as the code modifier. In our experiments, we use the original language model and GPT-4-0613 as code modifiers. For the sake of reproducibility, we choose the original LLM; To ensure the comparison focuses exclusively on modeling capacity, we choose GPT-4-0613 as the code modifier due to its exceptional coding proficiency and operational reliability, as well as its stability than other provisional release Language Models. See the comparison of the result from different code modifiers in Section 6.3, and prompts in Appendix  O.
@@ 6.1.2 Evaluation Metrics
Given the potential for minor discrepancies in numerical solutions, our protocol includes a comparison rule that accommodates for slight inaccuracies. Let $O$ denote the output from the LLM, and $A$ denote the standard answer. The comparison is defined by the following criteria  444 See the evaluation script in Appendix M
General Comparison Criterion Let $n$ denotes the decimal places of $A$ ($n=0$ if $A$ is integer). The Answer $O$ is correct if
$\left|\frac{O-A}{A}\right|<10^{-4}\text{ or }|O-A|<\min\{10^{-n},10^{-2}\}$ (1)
This metric is designed [^5] See the case study in Appendix D to account for the precision errors inherent in solvers and the specified significant figures in questions, ensuring that the LLM’s output is accurately evaluated against the standard answer with allowances for numerical precision and variability in computational responses.
@@ 6.2 Benchmarking Results
Table 3: Evaluation Results on the Mamo Benchmark. The subscript denotes improvement using the original LLM as a code modifier, which corrects syntax errors without changing the underlying logic, differentiating modeling from formatting errors (see Section 6.1.1). The "Overall" score represents the weighted average of correct rates, weighted by question count. The highest score in each category is marked in boldface; an underline signifies the top score with the LLM itself as code modifier. The score after using LLM as code modifier is computed as the sum of the original score and the value in the subscript.
Models ODE Optimization Overall (%)
First Second System Easy Complex
order (%) order (%) (%) LP (%) LP (%)
GPT-4o 67.86 +0.51 36.36 +0.91 40.00
+0.00
87.27
+0.15
22.75
+8.53
66.67 +1.73
GPT-4-turbo 64.80
+1.02
32.73
+0.91
40.00
+0.00
87.88 +0.00 23.22
+6.64
66.34
+1.41
GPT-4 59.18
+1.02
28.18
+0.91
40.00
+2.50
86.50
+0.46
21.08
+3.09
63.81
+1.12
GPT-3.5-turbo 16.33
+6.12
7.27
+4.55
10.00
+5.00
81.29
+3.53
9.48
+0.00
49.13
+3.48
Claude-3-Opus 55.61
+3.57
33.64
+1.81
45.00
+2.50
83.74
+0.31
9.48
+0.47
60.38
+1.08
Claude-3-Sonnet 52.04
+2.04
24.55
+0.90
27.50
+5.00
83.59
+0.61
24.64
+0.48
60.96
+0.99
Claude-3-Haiku 41.33
+9.18
23.64
+0.91
17.50
+2.50
78.53
+7.36
17.54
+1.42
54.84
+5.87
Gemini-1-pro 25.00
+8.16
13.64
+0.00
2.50
+2.50
78.83
+1.08
14.69
+0.00
50.45
+1.99
Gemini-1.5-pro 64.80
+1.02
30.00
+0.00
45.00 +5.00 85.28
+0.00
30.81 +5.21 66.09
+1.24
DeepSeek-v2 39.29
+1.02
20.00
+9.09
30.00
+0.00
86.50
+0.16
3.79
+0.00
56.49
+1.08
DeepSeek-coder-33B 28.57
+2.04
11.82
+0.00
12.50
+0.00
75.77
+1.22
18.96
+0.00
50.29
+0.99
DeepSeek-math-7b-base 3.57
+0.00
2.73
+0.00
2.50
+0.00
49.39
+1.22
6.64
+0.00
28.70
+0.66
DeepSeek-math-7b-rl 1.02
+0.51
1.82
+0.00
5.00
+0.00
20.55
+0.62
0.00
+0.47
11.58
+0.50
Llama-3-8b 0.51
+0.51
0.00
+0.00
0.00
+0.00
55.37
+0.46
6.64
+0.00
31.10
+0.33
Llama-3-70b 4.08
+0.51
1.82
+1.82
2.50
+0.00
66.10
+0.62
9.00
+0.00
38.13
+0.58
Qwen-1.5-72b-Chat 5.61
+6.12
4.55
+0.00
5.00
+7.50
65.03
+6.29
5.69
+0.00
37.55
+4.64
Qwen-1.5-110b-Chat 28.57
+4.08
6.36
+1.82
7.50
+0.00
65.18
+10.89
13.27
+2.37
43.93
+6.11
Mixtral-8x7b-instruct 23.47
+4.59
4.55
+1.81
10.00
+0.00
67.79
+2.15
9.95
+0.00
42.85
+2.06
Mixtral-8x22b-instruct 52.04
+1.53
11.82
+1.82
20.00
+0.00
81.90
+2.61
19.91
+1.42
57.82
+2.06
  • 1
    Here GPT-4-turbo denotes GPT-4-turbo-2024-04-09; GPT-4o denotes GPT-4o-2024-05-13; GPT-4 refers to GPT-4-0613; GPT-3.5-turbo denotes GPT-3.5-turbo-0125